the perimeter of a rectangle is 88 feet. the length of the rectangle is 4 feet longer than three times the width. find the dimensions of the rectangle.
Answer by checkley77(12844)

L=3W+4
2L+2W=88
2(3W+4)+2W=88
6W+8+2W=88
8W=88-8
8W=80
W=80/8
W=10 FT. FOR THE WIDTH
L=3*10+4=30+4=34 FT. FOR THE LENGTH.
PROOF:
2*34+2*10=88
68+20=88
88=88